OpenStudy (anonymous):

Simplify 11(2x + 3). can someone help me here?

OpenStudy (anonymous):

Ummmm my answers that I got are 11x + 5 22x + 3 22x + 33

OpenStudy (mathstudent55):

Multiply 11 by each term inside the parentheses.

OpenStudy (anonymous):

The answer is 22x + 33

OpenStudy (mathstudent55):

11(2x + 3) = 11*2x + 11*3 = 22x + 33.

OpenStudy (anonymous):

so it would be 11x 2 or 11x3?

OpenStudy (anonymous):

getting a note book.

OpenStudy (anonymous):

11*2x + 11*3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

so I have 11x2=22 and 11x3=33

OpenStudy (anonymous):

so my answer would be 22x+33?

OpenStudy (mathstudent55):

You have 11 * 2x = 22x and 11 * 3 = 33. When you put it together you end up with 22x + 33.

OpenStudy (anonymous):

thanks a lot! If I need more help may I ask u for the help?

OpenStudy (anonymous):

Simplify 8(3 - 2x). 24x - 16 16x - 24 24 - 16x I believe the right answer would be number 2.

OpenStudy (mathstudent55):

Multiply 8 * 3. What do you get?

OpenStudy (anonymous):

24

OpenStudy (mathstudent55):

Good. Notice you wrote correctly 24 (positive 24) and not -24. Notice also that B. has -24. B cannot be the answer.

OpenStudy (mathstudent55):

What is 8 * -2x?

OpenStudy (anonymous):

-16 or 16.

OpenStudy (mathstudent55):

Positive times negative is negative, so 8 * (-2x) = -16x. Now look in the choices, which choice has 24 and -16x?

OpenStudy (anonymous):

so it would be 24 - -16?

OpenStudy (anonymous):

ohhh got ya! thanks!

OpenStudy (anonymous):

24x - 16 16x - 24 24 - 16x

OpenStudy (anonymous):

those are my choices for this quiz. I think it would be the last one.

OpenStudy (mathstudent55):

Yes, it's the last one.

OpenStudy (anonymous):

now this one I was having troubles with too. Simplify 81.56 + n - 2.55. 79.10 - n 79.01 + n 81.56 - 2.55n

OpenStudy (anonymous):

I think I would do it 91.56 - 2.55?

OpenStudy (anonymous):

* 81.56

OpenStudy (mathstudent55):

Rewrite 81.56 + n - 2.55 as 81.56 - 2.55 + n and combine the numbers together.

OpenStudy (anonymous):

79.01 + n that's what I get.

OpenStudy (mathstudent55):

Pls post new problems as new posts.

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

okay will do